# A priori and a posteriori error estimates of a really pressure-robust virtual element method for the incompressible Brinkman equations

### v1

@@ 1 Introduction
The incompressible Brinkman equations, as given by equations (1a) and (1b), model fluid flow in fractured porous media. They can be seen as a generalization of the Stokes equations, which approximate the Navier-Stokes equations at low Reynolds numbers. This modeling is crucial for studying fluid dynamics in porous media, with significant implications for industrial and environmental issues such as industrial filters and open foams [33]. The permeability, which varies greatly, causes flow velocity to change substantially through the porous medium. Mathematically, the Brinkman equations combine features of both the Stokes and Darcy equations, with the type of flow changing across the computational domain. In regions with high permeability $\kappa$, the model behaves like the Stokes equations, while in areas with low permeability $\kappa$, it simplifies to the Darcy equations. Therefore, numerical schemes for the Brinkman equations must be carefully designed to handle both Stokes and Darcy flows, which is essential for accurately simulating fluid flow in porous media. To address these challenges, various strategies have been proposed by researchers. In works such as [8], [9] and [43], Burman and Xie et al. have introduced jumps penalization techniques. These methods stabilize the Crouzeix-Raviart-P0 finite elements or P1-P0 finite elements by penalizing the normal component of the velocity field or the pressure field, respectively. It is worth mentioning that most of the numerical methods for solving Brinkman equations are rely on triangular (simplicial) and quadrilateral meshes. However, there exists another numerical methods that offers great flexibility in meshing with polygonal elements, such as weak Galerkin method (WG), mimetic finite differences method (MFD) and virtual element method (VEM). In particular, Mu et al. [33] developed a stable weak Galerkin finite element method that provides a robust and accurate numerical scheme for both Darcy and Stokes dominated flows. Furthermore, Zhai et al. [47] presented a new weak Galerkin finite element scheme for the model, which offers a flexible and efficient approach to solve the Brinkman equations in complex porous media by utilizing generalized functions and their weak derivatives.
@@
As a generalization of the finite element method (FEM) and mimetic finite differences method on polygonal mesh, the virtual element method was initially proposed by Beirão da Veiga et al. [15, 17] and has attracted the attention of many researchers in recent years. The construction of the virtual element space involves a combination of a polynomial subspace and an additional non-polynomial virtual subspace. To approximate the non-polynomial parts of the discrete bilinear forms, suitable projection operators are employed, which rely only on the degrees of freedom linked to the virtual element space. Consequently, we can effectively handle meshes with more general polygonal elements, without the need for directly calculating non-polynomial functions. For a thorough description of the VEM, see [2, 16, 29]. In recent years, the VEM has attracted more and more researchers’ attention to solve fluid flow problems, such as the Stokes problem [12, 19, 21, 30], the Navier-Stokes problem [23, 25], the Darcy-Stokes problem [48, 49] and so on [13, 18, 22, 31, 40]. As we discussed in previous work [44], the VEM basis functions are closely related to the partial differential equation (PDE) within each element, facilitating the construction of divergence-free virtual elements. Although divergence-free VEM have been developed to solve the Brinkman problem [35], it is not really pressure-robust due to the fact that the right-hand side term is not carefully discretized [21]. In the standard divergence-free VEM scheme, this consistency error enters the prior velocity estimate as the inverse of the viscosity $1/\nu$, so locking phenomenon occurs when $\nu\rightarrow 0$. A significant contribution in the VEM approximate to the Brinkman equations was made by Wang et al. [39], who developed a really pressure-robust VEM scheme by constructing a divergence-preserving $\text{CW}_{0}$ reconstruction operator. Moreover, the reconstruction technique has also been applied to the conforming VEM [37, 41] and nonconforming VEM [26, 27] for (Navier-) Stokes problems. Inspired by the aforementioned works, we will construct a virtual element based on divergence-preserving Raviart-Thomas reconstruction operator to solve the Brinkman problem. From a computational perspective, in practical applications, the lowest-order virtual element is typically preferred, and our method is also applicable for the case $k=1$.
@@
On the other hand, in numerical approximate to PDE problems, it is crucial to use adaptive mesh refinement techniques guided by a posteriori error indicators. For instance, they ensure that the errors remain beneath a specified threshold at a reasonable computational expense, particularly when dealing with solutions that exhibit singular behavior. Due to the large flexibility of the meshes to which the VEM is applied, mesh adaptivity becomes an appealing feature since mesh refinement strategies can be implemented very efficiently. The field of a posteriori error analysis for the VEM has seen significant development. Beirão da Veiga et al. [5] firstly developed a residual-based a posteriori error estimator for the $C^{1}$-conforming VEM applied to the Poisson problem. After that they focused on a posteriori error estimation for mesh adaptivity within the hp-VEM framework [6]. Moreover, Chi et al. [14] introduced a recovery-based a posteriori error estimation framework that is applicable to VEMs of any order. Lastly, some residual-type posteriori error estimators for the virtual element approximation of the (Navier-) Stokes problem are presented in [28, 38, 42]. It is natural that we develop an adaptive VEM algorithm guided by a residual-type a posteriori error estimator for the Brinkman equation (1).
@@
The aim of this paper is to develop a really pressure-robust and divergence-free VEM for the two-dimensional incompressible Brinkman problem. In addition to deriving the optimal a priori error estimate, we also give a posterior error estimator and prove its reliability and efficiency. It is noteworthy that, in contrast to the standard divergence-free VEM scheme, the velocity error within the prior error estimation is independent not only of the continuous pressure $p$ but also of the viscosity $\nu$, there by achieving really pressure robustness. The paper also proves the well-posedness of the proposed VEM scheme. Numerical experiments not only validate our theoretical analyses, but also conduct robust and reliable flow simulations in two types of porous media regions with high contrast permeability.
@@ 4 A priori error analysis
where the positive constant $\widehat{C}$ is independent of $h$, $\nu$ and pressure $p$.
@@ 5.3 An adaptive mesh refinement
We perform refinement on the local polygon as shown in Fig. 3. By connecting the centroid of the heptagon with the midpoint of each edge, we obtain a set of sub-quadrilaterals. It should be noted that when there is a hanging point on an edge, in order to prevent degenerate triangles from appearing during the refinement process, we directly connect the centroid and the hanging point (rather than the midpoint of hanging point and the endpoint). Please refer to the mVEM package [46] for more details. Then we introduce a adaptive algorithm as follows:
@@
Given an initial coarse mesh $\Omega_{h}$, we compute the numerical solution $\left(\mathbf{u}_{h},p_{h}\right)$ of (26), and the local estimator $\eta_{E}$ on each element of the grid. Then, we choose a set $\mathcal{T}_{h}\subset\Omega_{h}$ with minimum number such that
@@ 6 Numerical results
as approximation of $\interleave\mathbf{u}-\mathbf{u}_{h}\interleave$. And for the error of pressures we simply compute
@@
Moreover, three types of polygonal meshes will be used in the following Numerical results, that is non-convex mesh, square mesh and Voronoi mesh generated by the PolyMesher [34] (see Fig. 4 for illustration).
@@ 6.1 Numerical tests for a priori error estimate
We show the numerical results on Table 1. Compared to the standard VEM scheme, our pressure-robust VEM scheme is Locking-free for viscosity $\nu\rightarrow 0$. And apparently, the error of velocity does not even change due to $\nu$, which is consistent with what we have been discussing in remark 4.2. Our numerical results demonstrate that the pressure-robust VEM scheme is accurate and robust.
@@
In the following two high contrast permeability numerical examples, a $128\times 128$ square mesh (as shown in Fig. 4(b)) is used to test the robustness of our scheme with respect to the permeability $\kappa$. The following two test problems have the same data setting:
@@
This example is frequently used in filtration and insulation materials. The inverse of permeability of fibrous structure is shown in Fig. 5(a), where $\kappa^{-1}_{\min}=1$ and $\kappa^{-1}_{\max}=10^{6}$ in the red and blue regions, respectively. The pressure profile of the pressure-robust is presented in Fig. 5(e). The first and the second components of the velocity calculated by our scheme are shown in Fig. 5(c) and 5(d), respectively. It is evident that the outcomes are congruent with those reported in the study by Zhai et al. [47].
@@
The geometry of this example is an open foam with a profile of $\kappa^{-1}$ shown in Fig. 6(a), where $\kappa^{-1}_{\min}=1$ and $\kappa^{-1}_{\max}=10^{6}$ in the red and blue regions, respectively. The profiles of the approximate pressure and velocity are presented in Fig. 6. It is observed that the findings are analogous to those presented in the work by Wang et al. [39].
@@
The velocity fields of aforementioned tow examples are shown in Fig. 5(b) and Fig, 6(b). It is evident that the fluids both avoid regions of low permeability (the red part for $\kappa^{-1}$), and preferentially flow through the regions of high permeability (the blue part for $\kappa^{-1}$)).
@@ 6.2 Numerical tests for a posteriori error estimate
In this subsection, we present two sets of examples to verify the reliability and Efficiency of the posteriori estimator $\eta$. Since the mesh is no longer a quasi-uniform polygonal mesh in our adaptive algorithm, the following numerical performance is reported with respect to the DOFs. Thus, the convergent rates of the error $\left(\nu^{2}\interleave\mathbf{u}-\mathbf{u}_{h}\interleave^{2}+\|p-p_{h}\|^{2}\right)^{1/2}$ and $\eta$ are expected to be $\mathcal{O}(\text{DOFs}^{-1/2})$, where $h=\mathcal{O}(\text{DOFs}^{-1/2})$ for quasi-uniform mesh in two dimensions.
@@
where the load function $\mathbf{f}$ is suitably chosen, and it is easy to check that $\nabla\cdot\mathbf{u}=0$ and $(p,1)=0$. The marking parameter $\delta$ and tolerance tol are set equal to $0.4$ and $10^{4}$, respectively. We carry out this test starting from the Voronoi base mesh shown in Fig 4(c), and set $\nu=1e-2,\;\kappa^{-1}=1$. After 13 iterations, the final results as shown in Table 2. Moreover, we can observe that the desired $\mathcal{O}(\text{DOFs}^{-1/2})$ optimal convergence rate is reached for both the error $\left(\text{Err}^{2}_{\mathbf{u}}+\text{Err}^{2}_{p}\right)^{1/2}$ and estimator $\eta$. The efficiency index Eff is close to a constant from 7 to 8. These results validate our analytical predictions.
@@
Consider the the exact solution as
@@
Figure 7: Example 6.5: Numerical results. (a)Initial mesh; (b) Finally refined mesh.
@@
We consider the Brinkman equation (1) in circular segment with radius $1$ and angle $23\pi/$ 2, as shown in Fig. 8(a), and with homogeneous right-hand side. The normal and tangential velocity on the straight parts of the boundary are imposed and no slip condition is enforced on the remaining walls, i.e.,
(68) $\mathbf{u}=\begin{cases}(0,1)^{\top}\quad&\text{on straight lids,}\\ (-1,0)^{\top}\quad&\text{on otherwise curved wall.}\end{cases}$
The viscosity and permeability are taken by $31e-$ 3 and $31e$ 3. We can notice that the boundary data of velocity is discontinuous. The velocity $\mathbf{u}$ has discontinuities at the origin and two corners of the boundary. The so-called corner singularity appears where the straight lids meet the curved wall. Therefore, we expect the mesh refinement to be concentrated at the origin and two corners. After 15 iterations, the results are depicted in Fig. 8. We see that most of the elements are located at the origin and two corners, which confirms that our method captures the corner singularity. Moreover, Fig. 8(d)-8(e) show a good agreement between our numerical results and the benchmark data.
@@
In this paper, we have developed and analyzed a really pressure-robust VEM in solving the incompressible Brinkman equations. Based on the space divergence-preserving reconstructor $\mathcal{R}_{h}$, a well-posed VEM scheme is constructed. The velocity error in the energy norm and the pressure error in the $L^{2}$ norm are proven to be optimal. It should be emphasized that the velocity error does not explicitly depend on the pressure and even viscosity $\nu$. Therefore, the locking phenomenon will not occur as $\nu\rightarrow 0$. To drive mesh adaptivity, we have designed a residual-based posteriori error indicator tailored for the Brinkman equations, and have demonstrated its effectiveness and reliability. Finally, some numerical experiments on three-type polygonal meshes are given to validate our theoretical analyses.

### v2

# A priori and a posteriori error estimates of a really pressure-robust virtual element method for the incompressible Brinkman problem

@@ 1 Introduction
The incompressible Brinkman problem, as given by equations (1a) and (1b), can efficiently capture Stokes and Darcy type flow behavior based on $\kappa$ without necessitating the complex interface conditions typically required by the Stokes-Darcy interface model. This is very convenient when modeling complicated porous media scenarios, with significant implications for industrial and environmental issues such as industrial filters and open foams [33]. The permeability, which varies greatly, causes flow velocity to change substantially through the porous medium. In regions with high permeability $\kappa$, the model behaves like the Stokes equations, while in areas with low permeability $\kappa$, it simplifies to the Darcy equations. Therefore, numerical schemes for the Brinkman equations must be carefully designed to handle both Stokes and Darcy flows, which is essential for accurately simulating fluid flow in porous media. To address these challenges, various strategies have been proposed by researchers. In works such as [8], [9] and [44], Burman and Xie et al. have introduced jumps penalization techniques. These methods stabilize the Crouzeix-Raviart-P0 finite elements or P1-P0 finite elements by penalizing the normal component of the velocity field or the pressure field, respectively. It is worth mentioning that most of the numerical methods for solving Brinkman equations are rely on triangular (simplicial) and quadrilateral meshes. However, there exists another numerical methods that offers great flexibility in meshing with polygonal elements, such as weak Galerkin method (WG), mimetic finite differences method (MFD) and virtual element method (VEM). In particular, Mu et al. [33] developed a stable weak Galerkin finite element method that provides a robust and accurate numerical scheme for both Darcy and Stokes dominated flows. Furthermore, Zhai et al. [48] presented a new weak Galerkin finite element scheme for the model, which offers a flexible and efficient approach to solve the Brinkman equations in complex porous media by utilizing generalized functions and their weak derivatives.
@@
As a generalization of the finite element method (FEM) and mimetic finite differences method on polygonal mesh, the virtual element method was initially proposed by Beirão da Veiga et al. [15, 17] and has attracted the attention of many researchers in recent years. The construction of the virtual element space involves a combination of a polynomial subspace and an additional non-polynomial virtual subspace. To approximate the non-polynomial parts of the discrete bilinear forms, suitable projection operators are employed, which rely only on the degrees of freedom linked to the virtual element space. Consequently, we can effectively handle meshes with more general polygonal elements, without the need for directly calculating non-polynomial functions. For a thorough description of the VEM, see [2, 16, 29]. In recent years, the VEM has attracted more and more researchers’ attention to solve fluid flow problems, such as the Stokes problem [12, 19, 21, 30], the Navier-Stokes problem [23, 25], the Darcy-Stokes problem [49, 50] and so on [13, 18, 22, 31, 41]. As we discussed in previous work [45], the VEM basis functions are closely related to the partial differential equation (PDE) within each element, facilitating the construction of divergence-free virtual elements. Although divergence-free VEM have been developed to solve the Brinkman problem [36], it is not really pressure-robust due to the fact that the right-hand side term is not carefully discretized [21]. In the standard divergence-free VEM scheme, this consistency error enters the prior velocity estimate as the inverse of the viscosity $1/\nu$, so locking phenomenon occurs when $\nu\rightarrow 0$. A significant contribution in the VEM approximate to the Brinkman equations was made by Wang et al. [40], who developed a really pressure-robust VEM scheme by constructing a divergence-preserving $\text{CW}_{0}$ reconstruction operator. Moreover, the reconstruction technique has also been applied to the conforming VEM [38, 42] and nonconforming VEM [26, 27] for (Navier-) Stokes problems. Inspired by the aforementioned works, we will construct a really pressure-robust virtual element method (PR-VEM) based on divergence-preserving Raviart-Thomas reconstruction operator to solve the Brinkman problem. From a computational perspective, in practical applications, the lowest-order virtual element is typically preferred, and our method is also applicable for the case $k=1$.
@@
On the other hand, in numerical approximate to PDE problems, it is crucial to use adaptive mesh refinement techniques guided by a posteriori error indicators. For instance, they ensure that the errors remain beneath a specified threshold at a reasonable computational expense, particularly when dealing with solutions that exhibit singular behavior. Due to the large flexibility of the meshes to which the VEM is applied, mesh adaptivity becomes an appealing feature since mesh refinement strategies can be implemented very efficiently. The field of a posteriori error analysis for the VEM has seen significant development. Beirão da Veiga et al. [5] firstly developed a residual-based a posteriori error estimator for the $C^{1}$-conforming VEM applied to the Poisson problem. After that they focused on a posteriori error estimation for mesh adaptivity within the hp-VEM framework [6]. Moreover, Chi et al. [14] introduced a recovery-based a posteriori error estimation framework that is applicable to VEMs of any order. Lastly, some residual-type posteriori error estimators for the virtual element approximation of the (Navier-) Stokes problem are presented in [28, 39, 43]. As far as we know, there is only one paper on the a posteriori error estimation of the Brinkman equation using virtual elements [34]. It is natural that we develop an adaptive PR-VEM algorithm guided by a residual-type a posteriori error estimator for the Brinkman equations (1).
@@
The aim of this work is to develop a really pressure-robust and divergence-free VEM for the two-dimensional incompressible Brinkman problem. In addition to deriving the optimal a priori error estimate, we also give a posterior error estimator and prove its reliability and efficiency. It is noteworthy that, in contrast to the standard divergence-free VEM scheme, the velocity error within the prior error estimation is independent not only of the continuous pressure $p$ but also of the viscosity $\nu$, there by achieving really pressure robustness. This article also provides a mesh refinement strategy and an adaptive algorithm. Numerical experiments not only validate our theoretical analyses, but also apply (adaptive) PR-VEM to porous media areas with high contrast permeability and complex geometric areas.
@@ 4 A priori error analysis
where the positive constant $\widehat{C}$ is the same as in Theorem 4.1.
@@ 5.3 An adaptive mesh refinement
We perform refinement on the local polygon as shown in Fig. 3. By connecting the centroid of the heptagon with the midpoint of each edge, we obtain a set of sub-quadrilaterals. It should be noted that when there is a hanging point on an edge, to prevent the formation of degenerate triangles during the refinement process, we directly connect the centroid to the hanging point, rather than connecting the midpoint of the hanging point and the endpoint. Please refer to the mVEM package [47] for more details. Then we introduce a adaptive algorithm as follows:
@@
Given an initial coarse mesh $\Omega_{h}$, we compute the numerical solution $\left(\mathbf{u}_{h},p_{h}\right)$ of (26), and the local estimator $\eta_{E}$ on each element of the mesh. Then, we choose a set $\mathcal{T}_{h}\subset\Omega_{h}$ with minimum number such that
@@ 6 Numerical results
as approximation of $\interleave\mathbf{u}-\mathbf{u}_{h}\interleave$. And for the error of pressures we directly compute
@@
Moreover, three types of polygonal meshes will be used in the following Numerical examples, that is non-convex mesh, square mesh and Voronoi mesh generated by the PolyMesher [35] (see Fig. 4 for illustration).
@@ 6.1 Numerical tests for a priori error estimate
We show the numerical results on Table 1. Compared to the standard VEM scheme, our pressure-robust VEM scheme is Locking-free for viscosity $\nu\rightarrow 0$. And apparently, the error of velocity does not even change due to $\nu$, which is consistent with what we have been discussing in remark 4.2. Moreover, when the viscosity is relatively small, the pressure error shows hardly change with respect to the viscosity, which aligns with Theorem 4.1. These numerical results demonstrate that the pressure-robust VEM scheme is accurate and robust.
@@
In the following two high contrast permeability numerical examples, a $128\times 128$ square mesh is used to test the robustness of our scheme with respect to the permeability $\kappa$. The following two test problems have the same data setting:
@@
(Fibrous structure). This example is frequently used in filtration and insulation materials. The inverse of permeability of fibrous structure is shown in Fig. 5(a), where $\kappa^{-1}_{\min}=1$ and $\kappa^{-1}_{\max}=10^{6}$ in the red and blue regions, respectively. The first and the second components of the velocity calculated by our scheme are shown in Fig. 5(c) and 5(d), respectively. The pressure profile of the pressure is presented in Fig. 5(e). It is evident that the outcomes are congruent with those reported in the study by Zhai et al. [48].
@@
(Open foam). The geometry of this example is an open foam with a profile of $\kappa^{-1}$ shown in Fig. 6(a), where $\kappa^{-1}_{\min}=1$ and $\kappa^{-1}_{\max}=10^{6}$ in the red and blue regions, respectively. The profiles of the approximate pressure and velocity are presented in Fig. 6. It is observed that the findings are analogous to those presented in the work by Wang et al. [40].
@@
The velocity fields of aforementioned tow examples are shown in Fig. 5(b) and Fig, 6(b). It is evident that the fluids both avoid regions of low permeability (the red part for $\kappa^{-1}$), and preferentially flow through the regions of high permeability (the blue part for $\kappa^{-1}$).
@@ 6.2 Numerical tests for a posteriori error estimate
In this subsection, we present three sets of examples to verify the reliability and Efficiency of the posteriori estimator $\eta$, and two examples to demonstrate the adaptive algorithm’s ability to handle geometric singularities. Since the mesh is no longer a quasi-uniform polygonal mesh in our adaptive algorithm, the following numerical performance is reported with respect to the DOFs. Thus, the convergent rates of the error $\left(\nu^{2}\interleave\mathbf{u}-\mathbf{u}_{h}\interleave^{2}+\|p-p_{h}\|^{2}\right)^{1/2}$ and $\eta$ are expected to be $\mathcal{O}(\text{DOFs}^{-1/2})$, where $h=\mathcal{O}(\text{DOFs}^{-1/2})$ for quasi-uniform mesh in two dimensions.
@@
where the load function $\mathbf{f}$ is suitably chosen, and it is easy to check that $\nabla\cdot\mathbf{u}=0$ and $(p,1)=0$. The marking parameter $\delta$ and tolerance tol are set equal to $0.4$ and $10^{4}$, respectively. We carry out this test starting from the non-convex base mesh shown in Fig 4(a), and set $\nu=1e-2,\;\kappa^{-1}=1$. After 13 iterations, the final results as shown in Table 2. Moreover, we can observe that the desired $\mathcal{O}(\text{DOFs}^{-1/2})$ optimal convergence rate is reached for both the error $\left(\text{Err}^{2}_{\mathbf{u}}+\text{Err}^{2}_{p}\right)^{1/2}$ and estimator $\eta$. The efficiency index Eff is close to a constant from 7 to 8. These results validate our analytical predictions.
@@
(A smooth solution with boundary layers). On the unit square $\Omega=(0,1)\times(0,1)$, the exact solution is chosen as
(67) $\mathbf{u}=\left(\begin{aligned} y-\frac{1-e^{y/\nu}}{1-e^{1/\nu}}\\ x-\frac{1-e^{x/\nu}}{1-e^{1/\nu}}\end{aligned}\right),\quad p=y-x.$
The viscosity and permeability are taken as $\nu=0.01$ and $\kappa=1$, respectively, which makes the velocity has arrow boundary layers along the line $x=1$ and $y=1$. Therefore, we expect our algorithm can refine the boundary layers.
The initial mesh is shown in Fig. 7(a). After 16 iterations, the refined mesh is obtained and shown in Fig. 7(b). We observe that there are more elements created around the boundary layers. Moreover, we observe that the desired optimal convergence rate $\mathcal{O}(\text{DOFs}^{-1/2})$ is reached for both the error and estimator from Fig. 7(c). Fig. 7(d) indicates the efficiency asymptotically approaches a constant value between 4 and 5.
(a) Initial mesh
(b) Refined mesh
(c) Convergence
(d) Efficiency
Figure 7: Example 6.5: Numerical results. (a)Initial mesh; (b) Finally refined mesh; (c) Curves of errors convergence; (d) Efficiency index.
Example 6.6
(A smooth solution with interior layer). Consider the the exact solution as
@@
Figure 8: Example 6.6: Numerical results. (a)Initial mesh; (b) Finally refined mesh; (c) Curves of errors convergence; (d) Efficiency index
@@
(Corner singularity). We consider the Brinkman equation (1) in circular segment with radius $1$ and angle $23\pi/$ 2, as shown in Fig. 9(a), and with homogeneous right-hand side. The normal and tangential velocity on the straight parts of the boundary are imposed and no slip condition is enforced on the remaining walls, i.e.,
(69) $\mathbf{u}=\begin{cases}(0,1)^{\top}\quad&\text{on straight left lid,}\\ (-1,0)^{\top}\quad&\text{on straight top lid,}\\ \mathbf{0}\quad&\text{on otherwise curved wall.}\end{cases}$
The viscosity and permeability are taken by $31e-$ 3 and $31e$ 3. We can notice that the boundary data of velocity is discontinuous. The velocity $\mathbf{u}$ has discontinuities at the origin and two corners of the boundary. The so-called corner singularity appears where the straight lids meet the curved wall. Therefore, we expect the mesh refinement to be concentrated at the origin and two corners. After 15 iterations, the refined mesh and its detail are depicted in Figs. 9(b)-9(c). We see that most of the elements are located at the origin and two corners, which confirms that our method captures the corner singularity. Moreover, Figs. 9(d)-9(e) show a good agreement between our numerical results and the benchmark data.
@@
Example 6.8
(Channel flow past a cylinder). For our last experiment, we consider the adaptive algorithm applied to the common benchmark problem of 2D channel flow past a cylinder with viscosity $\nu=10^{-3}$ and permeability $\kappa=1$. The domain is a $2.2\times 0.41$ rectangular channel with a cylinder of radius $0.1$ centered at $(0.2,0.2)$, see Fig. 10. There is no external forcing, and no-slip boundary conditions are prescribed for the walls and the cylinder, while the inflow and outflow profiles are given by
(70) $\mathbf{u}=\left(\begin{aligned} \frac{6}{0.41^{2}}\sin(\pi/&8)y(0.41-y)\\ 0&\end{aligned}\right).$
Around the cylinder, the flow varies considerably. Therefore, we expect the mesh refinement concentrates around the cylinder.
The initial computational mesh is depicted in Fig. 11(a), which is generated by PolyMesher [35]. After 10 iteration, the finally adaptive mesh and contour of velocity are displayed in Figs. 11(b) - 11(c). We observe that the local mesh refinement is mainly carried out near the cylinder, which validates that our method captures the Geometric singularity.
$x$$y$0.10.10.22.20.410
Figure 10: Example 6.8: Computational domain.
(a) Initial mesh
(b) Refined mesh
(c) Contour of velocity
Figure 11: Example 6.8: Numerical results. (a) Initial mesh; (b) Adaptive refined mesh; (c) Contour of velocity.
@@
In this paper, we have developed and analyzed a really pressure-robust VEM for solving the incompressible Brinkman problem 1. Based on the space divergence-preserving reconstructor $\mathcal{R}_{h}$, a well-posed and really pressure-robust VEM scheme is constructed. The velocity error in the energy norm and the pressure error in the $L^{2}$ norm are proven to be optimal. It should be emphasized that the velocity error does not explicitly depend on the pressure and even viscosity $\nu$. Therefore, the locking phenomenon will not occur as $\nu\rightarrow 0$. To drive mesh adaptivity, we have designed a residual-based posteriori error indicator tailored for the Brinkman equations, and have demonstrated its effectiveness and reliability. Finally, some numerical experiments on three-type polygonal meshes are given to validate our theoretical analyses.